# To Offload or Not To Offload: Model-driven Comparison of Edge-native and On-device Processing

### v1

@@ 4.4. Impact of Network Conditions
4.5. Impact of Request Rate
Figure 5. Latency comparison between on-device processing and offloading to A2 for MobileNetV2 with varying request rates under 10 Mbps and 20 Mbps networks.
Next, we demonstrate how the request rate $\lambda$ affects offloading decisions using the MobileNetV2 workload. In this experiment, both the TX2 and Orin Nano are set to their maximum performance modes. We evaluate edge offloading under 10 Mbps and 20 Mbps network bandwidth.
Figure 5 plots the end-to-end latency of each strategy as a function of request rate (RPS). When comparing Orin Nano with edge offloading to A2, we observe two distinct behaviors depending on the network bandwidth. At 10 Mbps, on-device processing using Orin Nano outperforms edge offloading across all request rates. However, under a 20 Mbps network, there exists a performance crossover point: On-device processing on Orin Nano provides lower latency under low load (RPS ¡ 60), but edge offloading becomes the faster option once the RPS exceeds 60. On the other hand, TX2 consistently exhibits the highest latency across all request rates due to its limited processing capability. Overall, the predicted end-to-end latencies of our models match closely with the observed latency, proving the accuracy of our models.
@@ 4.5. Impact of Request Rate
This result shows that while both the network queueing delays (for offloading) and local processing queueing delay (for on-device processing) grow with $\lambda$, the rate at which delays increase differs depending on the hardware and network configurations. Under 10 Mbps, the network queueing delay increases at a faster rate than the device queueing delay as $\lambda$ increases, favoring on-device processing. Conversely, under 20 Mbps, device queuing delays dominate at higher $\lambda$, shifting the advantage to offloading. These competing scaling effects validate the observation we discussed in Lemma 3.1.
@@
Key takeaway. As load increases, on-device processing becomes more advantageous under slower networks, while offloading provides lower latency under faster networks.
4.6. Impact of Multi-Tenancy
Figure 6. Latency comparison across execution strategies for InceptionV4 under a 100 Mbps network with varying number of co-located applications. Each application has RPS = 2.
@@ 4.6. Impact of Multi-Tenancy
We now experimentally validate our model’s ability to predict average response time in multi-tenant settings using Orin Nano as the device and A2 as the edge server. We vary the number of applications deployed on the edge, each serving InceptionV4 requests, with all models loaded in memory concurrently. Each application receives 2 requests per second from its corresponding client. We report the average offloading latency across all clients and compare it with on-device processing.
@@
Figure 6 compares the end-to-end latency of offloading with on-device processing as the number of co-located applications increases. When only a single application is deployed on the edge, offloading yields lower latency than on-device processing. As the edge server hosts more applications, the latency of offloading increases due to resource contention across applications. We observe a performance crossover point at 18 co-located applications, beyond which on-device processing becomes more efficient. Note that the on-device processing time remains constant regardless of the number of concurrent applications as devices operate independently. In all cases, our models closely predict the latency for both execution strategies.
@@
Key takeaway. Offloading latency increases with the number of co-located applications due to resource contention on the edge server.
5. Models in Action
@@
In this section, we present a resource manager that leverages our analytic models to adaptively decide between offloading and on-device processing. We outline its decision algorithm and demonstrate its effectiveness through two case studies that showcase how the resource manager adapts to network variability and workload interference in multi-tenant edge environments.
5.1. Model-Driven Adaptive Resource Management
@@
We implement a resource manager that utilizes our analytic models to adaptively switch between on-device and edge processing. The manager operates locally on the device and periodically collects runtime metrics such as network bandwidth, edge server load, and request arrival rate. For each epoch, it uses the models to estimate the average latency of both execution strategies and chooses the option with the lower predicted average latency to execute requests.
@@
Algorithm 1 outlines the decision-making process for selecting the execution strategy in the resource manager. At Lines 1-2, it predicts on-device average execution latency based on the current arrival rate and device processing speed. At Lines 3-6, it computes the predicted average latency for offloading to each edge server, considering the server load, service rates, and the device’s network conditions. Finally, at Lines 7-11, it executes the request using the best execution strategy based on these predictions.
@@ 2 $T_{dev}\leftarrow M/D/1(\lambda^{proc}_{dev},\mu_{dev}^{proc})+s_{dev}^{proc}$
3 $T_{net}^{req}\leftarrow M/M/1(\lambda_{dev},B/D_{req})+D_{req}/B_{dev}$
4 $T_{net}^{res}\leftarrow M/M/1(\lambda_{edge,E},B/D_{res})+D_{res}/B_{edge}$
@@ 10       $E^{*}\leftarrow\operatorname*{arg\,min}_{E}T_{edge,E}$
       return offload($E^{*}$) ;
        // Offload to server $E^{*}$
@@
Algorithm 1 Model-Driven Adaptive Offloading Algorithm
5.2. Case 1: Fluctuating Network Conditions
Figure 7. Observed latency under changing network bandwidth as the resource manager selects between offloading and on-device processing. Bandwidth is varied at t = 20, 40, and 60 seconds.
We first demonstrate how the resource manager adapts to changes in network conditions. We use the TX2 as the device, A2 as the edge server, and MobileNetV2 as the workload, with requests arriving at a fixed rate of 10 RPS. Initially, the device has a network bandwidth of 20 Mbps. To simulate network variation, we reduce the bandwidth to 10 Mbps at t = 20, further reduce it to 2 Mbps at t = 40, and then increase it back to 20 Mbps at t = 60.
@@ 5.2. Case 1: Fluctuating Network Conditions
Key takeaway. Our resource manager dynamically adapts to changing network conditions by switching between execution strategies.
5.3. Case 2: Multi-tenant Edge Servers
@@
Figure 8. Observed latency under varying edge server workloads. At t = 80, the request rate of edge server 1 is increased from 10 RPS to 50 RPS. At t = 160, the request rate of edge server 2 is also increased to 50 RPS.
@@
Next, we demonstrate how our resource manager adapts to workload changes across multiple multi-tenant edge servers. In this experiment, two edge servers, $E_{1}$ and $E_{2}$, each equipped with an A2 GPU, are hosting separate instances of a YOLOv8n application. Initially, the request rates are set to $\lambda_{E_{1}}=10$ and $\lambda_{E_{2}}=30$, while a TX2 running the resource manager has a request rate of $\lambda_{dev}=10$. At t = 80, we add an additional device offloading request to $E_{1}$ to increase $\lambda_{E_{1}}$ to 50. Similarly, at t = 160, we increase $\lambda_{E_{2}}$ to 50.
@@
Figure 8 shows the observed latencies alongside the predicted average latencies for offloading to $E_{1}$, $E_{2}$, and on-device processing. Initially, from t = 0 to t = 80, the predicted latency for offloading to $E_{1}$ is the lowest, therefore the resource manager offloads requests to $E_{1}$. However, at t = 80, as the workload on $E_{1}$ increases, its predicted latency rises. In response, the resource manager adapts by offloading to the less loaded server $E_{2}$. At t = 160, as both servers reach a request rate of 50 RPS, the predicted latency for on-device processing becomes the lowest, and it shifts to on-device processing. This sequence demonstrates the resource manager’s ability to dynamically adapt to fluctuations in edge server workload.
@@
Key takeaway. Our resource manager dynamically adapts to workload changes in multi-tenant edge servers to maintain low latency.
6. Related Work
@@
Edge offloading. Edge offloading has become a widely adopted strategy in many applications to enable resource-constrained devices to perform computationally intensive tasks by leveraging edge infrastructure (Davis et al., 2004; Jang et al., 2018; Lai et al., 2017; Li et al., 2024; Ogden et al., 2021; Ren et al., 2019; Mei et al., 2012; Liu et al., 2018; Chen et al., 2017; Wang et al., 2018; Zhang et al., 2024a; Tkachenko et al., 2023). For instance, Furion (Lai et al., 2017) offloads computation-intensive components of mobile AR pipelines to the edge. Jang et al.(Jang et al., 2018) propose an edge camera virtualization architecture to offload video analytics workloads. Li et al.(Li et al., 2024) introduce a hybrid chaotic evolutionary algorithm to dynamically determine the optimal offloading location of DNN layers under energy constraints. In contrast, our work uses a model-driven approach to reexamine whether offloading is advantageous in the era of accelerators. We extend our analysis to capture the combined effects of hardware characteristics, workload and network dynamics, and multi-tenancy in edge servers.
@@
Model-aware placement and resource allocation. Queueing models have been extensively used to predict the performance of traditional CPU-bound workloads (Jiang et al., 2021; Wang et al., 2024a; Tantawi and Towsley, 1985; Ambati et al., 2020; Gandhi and Suresh, 2019; Shen et al., 2011; Urgaonkar et al., 2005). Recently, there has been work focusing on modeling the performance of accelerators. For example, Reference (Merck et al., 2019) models the latency of GPU batching in DNN applications. Liang et al. (Liang et al., 2023) model the performance of executing inference requests on various accelerators (e.g., TPU, GPU) and propose an online knapsack algorithm that uses the models for DNN model placement. In contrast, our models incorporate not only device hardware characteristics but also network delays and device-to-edge performance differences, enabling direct comparisons between on-device and edge offloading strategies.
Collaborative inference. There is a growing body of work on collaborative inference, where the computation is split across multiple nodes in the network to reduce end-to-end latency (Almeida et al., 2022; Banitalebi-Dehkordi et al., 2021; Gao et al., 2022; Hu et al., 2019; Huang et al., 2020; Kang et al., 2017; Laskaridis et al., 2020; Li et al., 2018; Merck et al., 2019; Ng et al., 2024; Ogden and Guo, 2023; Tuli et al., 2023; Yang et al., 2022; Zhang et al., 2020, 2024b). Neurosurgeon (Kang et al., 2017) proposes regression models for layer-wise execution times prediction for DNN partitioning across network components. SPINN (Laskaridis et al., 2020) integrates early-exit networks with model partitioning across device and cloud, enabling graceful degradation under network instability. In this work, we focus on offloading entire inference requests and do not consider partitioned execution.
7. Conclusions
@@
In this paper, we presented analytic models to characterize the latency behavior of both on-device processing and edge offloading using hardware accelerators. We validated the accuracy of our model across a range of hardware, workloads, and network configurations. We developed a resource manager that leverages these predictions to dynamically adapt to changing network conditions and multi-tenant edge server workloads. As future work, we plan to extend our models to incorporate collaborative processing, i.e., performing a portion of request computation locally before offloading.

### v2

@@ IV-E Impact of Network Conditions
IV-F Impact of Collaborative Processing
To demonstrate our models’ ability to capture collaborative (i.e., split) processing behavior, we consider an example using MobilNetV2 and use the Orin Nano as the device and A2 as the edge under a 50 Mbps network. We evaluate three split points (SPs) where the model is progressively offloaded to the edge and compare them to full on-device execution. For each split point, the initial portion of the neural network is executed locally on the device, while the resulting intermediate output and remaining computation are offloaded to the edge. Figure 6a plots the observed latencies of each strategy alongside model predictions. To compute the end-to-end latency, we use the combined analytic models for device and edge processing as discussed in Section III-E. As shown, the latency increases with the amount of compute offloaded. This is because the intermediate results of later layers grow in size, adding to the network transfer overhead. As seen, for MobileNetV2 and the chosen network configuration, local processing is faster than split processing. Across all split points, the model-predicted latencies closely align with observed values. Note that other models will exhibit other types of behaviors with split processing, which we omit due to space constraints.
Key takeaway. Our models can be extended to capture both compute and network delays in device-edge collaborative processing.
IV-G Impact of Request Rate
Figure 6b shows how the request rate $\lambda$ affects offloading decisions using the MobileNetV2 workload under 10 Mbps and 20 Mbps network bandwidth, with both the TX2 and Orin Nano set to their maximum performance modes. We observe two distinct behaviors depending on network bandwidth. At 10 Mbps, on-device processing using Orin Nano outperforms offloading across all RPS. However, at 20 Mbps, a performance crossover occurs: on-device processing is faster at low load (RPS $\leq$ 60), while offloading becomes more efficient beyond that point. In contrast, processing on TX2 consistently exhibits higher latency across all RPS due to its limited performance. In all cases, as the system approaches saturation at high RPS, queue lengths show increased variability from the Poisson arrival process, leading to greater fluctuations in queuing delays and response times. Overall, the predicted end-to-end latencies of our models match closely with the observed latency within confidence intervals, demonstrating the accuracy of our models.
@@ IV-G Impact of Request Rate
This result validates the competing scaling effects of $\lambda$ discussed in Lemma III.1. Although both the network queuing delays (for offloading) and local processing queuing delay (for on-device processing) grow with $\lambda$, the rate at which the delays increase differs depending on the hardware and network configurations. At 10 Mbps, the network queuing delay grows faster than the device queuing delay as $\lambda$ increases, favoring on-device processing. Conversely, at 20 Mbps, device queuing delays become dominant at higher $\lambda$, shifting the advantage toward offloading.
@@
Key takeaway. As load increases, on-device processing becomes more advantageous under slower networks, while offloading provides lower latency on faster networks.
IV-H Impact of Multi-Tenancy
@@
We now experimentally validate the accuracy of our models in multi-tenant settings, using Orin Nano as the device and A2 as the edge server. We vary the number of applications deployed on the edge, each serving InceptionV4 requests, with all models loaded in memory concurrently. Each application receives 2 RPS from its corresponding client. We report the average offloading latency and compare it with on-device processing.
@@
Figure 6c compares the average latency of offloading with on-device processing as the number of co-located applications increases. We observe that when only a single application is deployed on the edge, offloading yields lower latency than on-device processing. However, as the number of co-located applications increases, offloading latency also rises. This increase is caused by resource contention among applications because the A2 GPU does not support isolation. As a result, there is a crossover point at 18 applications, beyond which on-device processing becomes more efficient. Note that the on-device processing latency remains stable regardless of the number of concurrent applications as each device operates independently. In all cases, our models closely predict the latency for both execution strategies.
@@
Key takeaway. Offloading latency increases with the number of co-located applications due to resource contention on edge servers lacking GPU isolation.
V Models in Action
@@
In this section, we present a resource manager that leverages our analytic models to adaptively switch between offloading and on-device processing. We describe its decision algorithm and demonstrate its effectiveness through two case studies showcasing how the models enable adaptation to network variability and multi-tenant interference at the edge.
V-A Model-Driven Adaptive Resource Management
@@
The resource manager runs locally on the device and periodically collects runtime metrics, including network bandwidth, edge server load, and request arrival rate. At each epoch, it inputs these metrics into our models to estimate the average latency of both execution strategies and selects the one with the lower predicted latency to execute requests.
@@
Algorithm 1 outlines the decision-making process for selecting the execution strategy. Lines 1-2 predict the on-device average execution latency based on the current arrival rate and the estimated service time on the device. Lines 3-6 compute the predicted average latency for offloading to each edge server, considering server load, service rates, and network conditions. Finally, Lines 7-11 execute the request using the strategy with the lower predicted latency.
@@ 2 $T_{dev}\leftarrow M/D/1(\lambda^{proc}_{dev},\mu_{dev}^{proc})+s_{dev}^{proc}$
3 $T_{net}^{req}\leftarrow M/M/1(\lambda_{dev},B/D_{req})+D_{req}/B$
4 $T_{net}^{res}\leftarrow M/M/1(\lambda_{edge,E},B/D_{res})+D_{res}/B$
@@ 10       $E^{*}\leftarrow\operatorname*{arg\,min}_{E}T_{edge,E}$
       return offload($E^{*}$) // Offload to $E^{*}$
@@
Algorithm 1 Model-Driven Adaptive Offloading
V-B Case 1: Fluctuating Network Conditions
Figure 7: Observed latency under changing network bandwidth as the resource manager selects between offloading and on-device processing. Bandwidth is varied at t = 20, 40, and 60 seconds.
We first demonstrate how the resource manager leverages our models to adapt to changing network conditions. We use TX2 as the device, A2 as the edge, with MobileNetV2 requests arriving at a fixed rate of 10 RPS. Initially, the device has a network bandwidth of 20 Mbps to represent a stable 5G connection. To simulate real-world network variation, we reduce the bandwidth to 10 Mbps at t = 20 to reflect moderate signal degradation, further reduce it to 2 Mbps at t = 40 to emulate severe network congestion or poor coverage, and then restore it to 20 Mbps at t = 60 to represent signal recovery.
@@ V-B Case 1: Fluctuating Network Conditions
Key takeaway. Our models enable the resource manager to dynamically adapt to changing network conditions by switching between execution strategies.
V-C Case 2: Multi-tenant Edge Servers
@@
Figure 8: Observed latency under varying edge server workloads. At t = 80, the request rate of edge server 1 is increased from 10 RPS to 50 RPS. At t = 160, the request rate of edge server 2 is also increased to 50 RPS.
@@
Our second case study demonstrates how the resource manager leverages our models to adapt to workload changes across multiple multi-tenant edge servers. In this experiment, two edge servers, $E_{1}$ and $E_{2}$, each equipped with an A2 GPU, are hosting separate instances of a YOLOv8n application. Initially, the request rates of each server are set to $\lambda_{E_{1}}=10$ and $\lambda_{E_{2}}=30$, while the TX2 running the resource manager has a request rate of $\lambda_{dev}=10$. At t = 80, we add an additional device offloading request to $E_{1}$ to increase $\lambda_{E_{1}}$ to 50. Similarly, at t = 160, we increase $\lambda_{E_{2}}$ to 50.
@@
Figure 8 shows the observed latencies alongside the predicted average latencies for offloading to $E_{1}$, $E_{2}$, and on-device processing. Initially, from t = 0 to t = 80, the predicted latency for offloading to $E_{1}$ is the lowest, therefore the resource manager offloads requests to $E_{1}$. However, at t = 80, as the workload on $E_{1}$ increases, its predicted latency rises. In response, the resource manager adapts by offloading to the less loaded server $E_{2}$. At t = 160, as both servers reach a request rate of 50 RPS, the predicted latency for on-device processing becomes the lowest, and the resource manager shifts to on-device processing. This sequence highlights how our model enables the resource manager to dynamically adapt to fluctuations in edge server workload.
@@
Key takeaway. Our models enable the resource manager to dynamically adapt to workload changes in multi-tenant edge servers to maintain low latency.
VI Related Work
@@
Edge offloading. Edge offloading is a widely adopted strategy in many applications to enable resource-constrained devices to perform computationally intensive tasks by leveraging edge infrastructure [5, 8, 43, 44, 45, 6, 46, 47, 48, 49, 50, 51, 52]. For example, Furion [43] offloads computation-intensive components of mobile AR pipelines to the edge. Li et al. [44] introduce a hybrid chaotic evolutionary algorithm to dynamically determine the optimal offloading location of DNN layers under energy constraints. In contrast, our work uses a model-driven approach to reexamine whether offloading is advantageous in the era of accelerators. Our models apply to a wide range of workloads and generalizes to device-edge collaborative processing as shown in the evaluation.
@@
Model-aware placement and resource allocation. Queuing models have been widely used to predict the performance of traditional CPU-bound workloads [53, 3, 54, 55, 56, 57, 20]. More recently, there has been growing interest in modeling the performance of hardware accelerators. For example, Merck et al. [58] model the latency of GPU batching in DNN inference workloads. Liang et al. [29] develop performance models for inference on accelerators (e.g., GPU, TPU) and use them to guide DNN placement via an online knapsack algorithm. In contrast, our models are applicable not only to inference workloads but also to other accelerator-driven applications as shown in Section IV-D. They jointly capture device hardware characteristics, network delays, and device-to-edge performance differences, enabling direct comparison between on-device processing and edge offloading.
Modeling distributed edge workloads. Several studies have developed analytic models for distributed device–edge–cloud environments. Raei et al. [59] analyze the impact of wireless signal strength on request rejection probability and mean response delay. FogQN [60] uses multi-class queuing networks to model fog and cloud computing to determine the optimal fraction of data processing executed at the cloud versus at fog servers. Loghin et al. [61] model MapReduce applications in hybrid edge–cloud environments. However, these works abstract away accelerator-specific queuing behavior and do not consider the effects of multi-tenancy at the edge. Our work complements existing efforts by addressing these missing dimensions through a unified queuing-based modeling framework that informs resource management decisions.
VII Conclusions
@@
This paper presented analytic models to characterize the latency behavior of both on-device processing and edge offloading using hardware accelerators. We validated the accuracy of our model across a range of hardware, workloads, and network configurations. We developed a resource manager to demonstrate how model predictions can be used to dynamically adapt the execution strategy in response to changing network conditions and multi-tenant edge server workloads. As future work, we plan to extend the framework to model continuous batching in LLM inference and to support multi-stage pipelines or DAGs with dependencies across stages.